A horse canters away from its trainer in a straight line, moving 106 m away in 14.0 s . It then turns abruptly and gallops halfway back in 4.6 s .
1
Expert's answer
2020-01-27T09:33:21-0500

Total distance=106m

Time taken to covers this distance =14 sec

So velocity of the horse ="\dfrac{106}{14}=7.57 m/sec"

In the returning, the distance covered "=\dfrac{106}{2}=53m"

times=4.6 sec

So, returning speed ="\dfrac{53}{4.6}=11.52 m/sec"